shift summary for day shift 11.13.99

Jim Graham (jagraham@hep.uchicago.edu)
Sat, 13 Nov 1999 16:33:14 -0600

Shift Summary for Day shift, 11.13.99

On Shift: JAG, SP, with guest appearance by EB, ST and RJT
(this summary was ghostwritten by RJT)

Hours of beam delivered at all intensities: 8.25
Hours of beam delivered at nominal intensity: 8.00
Hours of E799 data (special runs not included): 7.25
Hours of special runs (laser scans, muon runs...): 0.00

Accounting of lost time:
SCSI error on KTeV4 (see log xxii p215) Run 14957 aborted
No success starting runs 14958-14960

========================================================================
Runs ended this shift: 14957

Run 14957 - 13.7M events - BEAM_799

========================================================================
Accesses during shift: None

========================================================================
Overview of shift:
Smooth E799 data taking (modulo SCSI errors noted above)
Accelerator ran exceptionally well this shift. Run 14597
began with ~7.5e12 but quickly went to 9e12 around 0730
remained there for the entire shift.

========================================================================
Description of Shift:

run spill time comment
------------------------------------------------------------------------

14957 7:00 Steven Pordes and Ed Blucher arrive to take over
shift. Steven does detailed survey of kumacs. Minor
problems in cosmic counters, and timing shifts
in RF, VV' and muon counters. Also see evidence
for 'mad latch' type behavior in VV' due to timing
shift. Since we have TDCs, don't fret about it.
8:20 Jim arrives.
9:25 mount onsplit tape UPM027
mount for onsplit tape UPF034 fails, and drive
on ktev4 refuses to relinquish the tape
11:49 run aborts
14958 -- Can't start new runs, because RC cannot write tape
14960 label to drive4 on ktev4. Eventually Sydney tracks
down the problem in the log file, and resets the
8 bay containing the two delinquent drives. This
solves both problems
14961 12:55 start run.
13:40 Rick T valiantly comes in to allow Jim to go the
832 meeting.
14961 79 14:40 Pipeline Errors; reset pipeline & run KFREND.

========================================================================
Follow up Notes on Previous Shift

========================================================================
Items to Follow up on Next Shift
- Monitor timing L1 timing shift
- Monitor CPU usage on KTeV 4; process 37 on ktev4 appears not to be
processing data.